OpenStudy (anonymous):

PLZ HELP ME!! -27 DIVIDED BY -9

OpenStudy (anonymous):

-9 * 3 = -27 -27 divided by -9 is 3

OpenStudy (anonymous):

-27/-9 negatives cancel, 27/9 3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

+3

OpenStudy (anonymous):

cuz -9*3=-27

OpenStudy (anonymous):

i could have sworn i just answered this, while fiddlaround watched